# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

### v1

@@ Abstract
Image captioning is fundamental for applications like video instruction systems and exploration robots, yet deploying such models on local devices is challenging due to the high computational demands of multimodal large language models (MLLMs). To address this, we first explore lightweight captioning by implementing a specialist based on a 125M-parameter language model, 56 times smaller than LLaMA-7B, and evaluating its performance on both single-sentence and detailed captioning tasks. Surprisingly, we find that our model can achieve performance comparable to large multimodal generalists, suggesting its potential to serve as a strong visual specialist for on-device applications. While promising, our model also exhibits a limitation: like other MLLMs, it suffers from visual blindness, occasionally resulting in semantic captioning errors. We carry out toy experiments and investigate the underlying causes, where we observe that the problems arise from ineffective attention mechanisms and limited visual representations. To alleviate them, we develop a novel captioning framework, Sharp-Eyed Refinement, which enhances caption quality through improved visual grounding. At its core, our DeepLens extracts detailed visual representations by concentrating on informative regions identified during the initial glance. Our experiments confirm both the advantages of our specialist over prior small captioning models and large generalists and the effectiveness of our framework.
Figure 1: (a) Image captioning serves as the cornerstone of diverse applications, such as robots for assistive navigation. (b) The natural process of image description is illustrated: humans first take in the overall scene, then glance at specific regions to notice finer details. (c) Our Sharp-Eyed Refinement framework mimics the human tendency to take a second, more focused glance, allowing the captioning specialist to revise and improve the initial descriptions.
@@ 1 Introduction
Image captioning has rapidly advanced with the emergence of MLLMs [43; 37], leveraging the impressive language processing capabilities of LLMs [94]. This captioning technology can play a crucial role in various applications, as shown in Fig. 1 (a). For example, video instruction systems utilize frame-wise caption generation for temporal understanding [76; 83; 78], while exploration robots construct graph-structured scene descriptions to navigate complex environments [82; 28].
@@
Despite these advantages, deploying such technologies on local devices remains a significant challenge due to the large computational demands of MLLMs [29]. Hence, we explore lightweight captioning by implementing a captioning specialist using OPT-125M [92], a model 56 times smaller than LLaMA-7B [67] in the LLaVA framework [43]. Evaluating it on both single-sentence and detailed captioning tasks, we found that our lightweight model can achieve impressive performance. In fact, it not only surpasses recently introduced small-scale captioning models [58] but also achieves performance comparable to large-scale multimodal generalists, such as LLaVA-7B [43] and Instruct-BLIP [15]. These results are particularly notable considering the limited language modeling capacity of OPT-125M, and suggest that the complex reasoning capabilities of LLMs may be less critical for image captioning tasks centered on factual information.
@@
Despite these remarkable results, our lightweight captioner is not without limitations. It occasionally generates inaccurate captions, an issue widely studied in MLLMs, such as LLM’s hallucination [5; 2] and the blindness of vision encoders [65; 66]. Since the majority of our specialist’s parameters are dedicated to the visual encoder, we focus on visual blindness as a potential root cause. This issue implies that since the vision encoder manipulates visual representation while perceiving the image with its eyes wide shut [66], the LLM often struggles to generate accurate outputs.
@@
We take a step towards analyzing the limitations of current captioning models. First, we inspect the model’s attention during a single pass captioning and find that the model struggled to focus on critical regions of the image, instead spreading its attention across the entire image. Moreover, we observe that visual representations from a CLIP encoder [56] lack sufficient detail for accurate caption generation. Specifically, when reconstructing images using these features, significant deviations from the original images occurred. These analyses highlight the need for more focused and detailed visual representations to provide the LM with clear detection information.
@@
Building on these insights, we propose a novel captioning framework, Sharp-Eyed Refinement, as illustrated in Fig. 1. Our method mirrors the human process of image description, where an initial broad understanding of the image is followed by a more focused and detailed analysis. In our framework, our model generates a rough caption, then refines its description. At the heart of our framework is our multimodal connector, DeepLens, which facilitates efficient re-examination of both the image and the initial caption during refinement.
@@
Our contributions are summarized as follows: (i) We investigate the practically important yet underexplored topic of lightweight captioning models (Sec. 2), and demonstrate their potential as strong visual specialists (Sec. 3). (ii) We suggest embodying the human mechanism of ‘one more glance with sharp eyes’ and introduce a novel refinement framework (Sec. 4). (iii) We present DeepLens, which extracts relevant visual features guided by the initial caption to enable focused refinement (Sec. 5). (iv) We conduct extensive experiments to validate the effectiveness and efficiency of our proposed framework and module. (Sec. 6). It is our hope that these efforts will inspire further development of efficient captioning models deployable on resource-constrained devices.
@@ 2 Motivation & Scope
Image Captioning as Foundational Technology
Image captioning is the process of transforming visual information from an image into a natural language description [51; 24]. More than just a standalone task, this technology serves as a fundamental component for a wide range of applications, as outlined in Fig. 1 (a). To process video input in chatbot systems [61; 77; 95; 83], image captioning is employed to generate descriptions of multiple frames, which are then integrated into LLMs as prompts to facilitate instruction following. Similarly, exploration robots [82; 28] designed for disaster environments, as well as assistive robots for the visually impaired, utilize captioning to encode observed scenes into structured graph representations, supporting navigation and user interactions. In this work, we focus on image captioning as a crucial enabler for these applications.
@@
Recent studies aiming to implement such applications commonly utilize open-source MLLMs [76] or cloud-based APIs (e.g., OpenAI API) [82] as image captioners. From a practical application perspective, however, these approaches present significant challenges, as (1) open-source MLLMs require substantial computational resources that exceed the capabilities of edge devices like smartphones [7; 63], while (2) cloud-based APIs depend on a stable internet connection, which is likely to be unavailable in disaster situations. Tab. 1 presents the number of parameters used by recent MLLMs and their NVIDIA GPU memory requirements, assuming FP16 precision. Given that these applications require repetitive captioning process across multiple scenes [90], the feasibility of current models for real-world deployment remains uncertain. This resource-demand mismatch highlights the need for lightweight captioning models.
@@ 3 Exploring Lightweight Captioning
3.1 Model Construction & Experimental Setup
Model construction. To construct a lightweight captioning model, we initially focused on minimizing reliance on LLMs, as they contribute to the majority of computational overhead in MLLMs (e.g., 96% in LLaVA-7B [43] is attributed to LLaMA [67]). To this end, we simply build our captioning specialist by replacing the LLaMA-7B used in LLaVA construction with OPT-125M [92], a language model 56 times smaller. Released in 2022, OPT-125M is significantly compact unlike modern small-scale language models exceeding 500M parameters.
Implementation details. We leverage the publicly available LLaVA-1.5 [42] code. Other than this modification, all training strategies follow the original code, including batch size, learning rate, and updating both the language model and the connector during fine-tuning while freezing the vision encoder. More details are available in Sec. H.1 and at https://github.com/junha1125/Lightweight-Captioner. All our experiments are conducted using two NVIDIA A6000 GPUs.
Datasets and Evaluations. The multimodal connector is first pretrained on the Caption Concept-balanced 558K [42], after which the model is fine-tuned on task-specific datasets such as COCO [14], DCI [70], or ShareGPT4V [11], as introduced later. We consider traditional metrics such as BLEU [55], CIDEr [71], and BERTScore [93], as well as MLLM-as-a-Judge [9; 8] with GPT-4o-mini.
@@ 3.2 Results on Single Sentence Captioning
MS COCO Captions [14], the most widely used captioning benchmark, consists of 113k images and 5 captions per image. One caption typically contains a single sentence of about 10 words. The results in Tab. 2 show that our model outperforms previous small captioning works [25; 72] that utilize fewer than 1B parameters. Specifically, our model achieves CIDEr scores that are 6.9 higher than SmallCap [58]. Considering that our model employs a small LM and does not involve any specialized methods, these results may represent surprisingly strong performance. Furthermore, we compare our specialist model against large multimodal generalists [13; 3] that leverage LLMs and are trained on proprietary instruction-following datasets [43]. Despite requiring significantly fewer computational resources, our specialist model achieves comparable performance to these larger counterparts. A discussion of these results is provided in Tab. 14 and Sec. D.1.
MS COCO [14]
model venue # data # params B@4 [55] MET [16] CIDEr [71] BERT [93] CLAIR [8] GPT [9]
G InstructBLIP [15] NeurIPS23 130M> 8.2B 38.0 29.4 127.8 69.1 - -
Unified-IOXL [46] ICLR23 130M 7.3B 37.0 29.5 123.6 68.2 - -
Shikra [10] arXiv23 - 7.2B - - 117.5 - - -
@@
ShareGPT4V [11] & DCI [70]
model venue # data # params B@4 [55] CIDEr [71] BERT [93] CAPT [18] CLAIR [8] GPT [9]
G InstructBLIP [15] NeurIPS23 130M> 8.2B - 28.2 - 41.7 - -
@@
SmallCap* [58] CVPR23 0.5M 450M 14.5 20.1 28.9 23.3 - -
@@
GLaMM [60]
model venue # data # params B@4 [55] CIDEr [71] BERT [93] CAPT [18] CLAIR [8] GPT [9]
[…]
Table 2: We evaluate our specialist and existing captioning models. Here, G represents a large multimodal generalist trained with instruction data to handle multiple tasks, while "S" denotes the small-scale captioning specialist trained on a dedicated captioning training dataset. * indicates models that are fine-tuned in this study, as they were not trained for detailed captioning tasks.
@@ 3.3 Results on Detailed Captioning
Images can be described with more detailed content beyond single-sentence descriptions [21]. Therefore, we continue our study by examining our specialist’s capabilities on more challenging detailed captioning tasks. For this experiment, we perform fine-tuning with the DCI [70] and ShareGPT4V [11] datasets, containing 100K and 7.4K training images, respectively. We then evaluate our model on the test sets of the above datasets. We also use the GLaMM [60] dataset, which contains 100K training and 6K testing images. Dataset details are included in Sec. F.2.
Figure 3: Generation examples from the original OPT-125M. Despite such limited capabilities, lightweight language models can perform factual image captioning efficiently.
@@
We initially anticipated that our specialist would struggle with this challenging task, given that OPT-125M possesses limited language processing capabilities, presumably due to its small parameter count, as shown in Tab. 3. However, the results in Tab. 2 and Fig. 4 reveal that our model performs surprisingly well on detailed captioning tasks, contrary to our expectations. We note that we keep the vision encoder frozen while only updating OPT-125M. Thus, detailed caption generation relies entirely on OPT-125M, with the vision encoder providing only visual cues.
Figure 4: Qualitative results on detailed captioning. Our specialist, despite its small size, can generate effective image descriptions. More examples are provided in Sec. E.
@@
These results underscore the efficiency of a lightweight LM for captioning tasks. Although MLLMs rely on the complex reasoning abilities of LLMs for tasks such as VQA and instruction following, our experiments indicate that captioning, primarily involving the enumeration of factual objects, does not significantly require such capabilities of LLM. This suggests that a lightweight captioning model can serve as an efficient and practical alternative for captioning-based applications.
Note 1:
@@
Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
@@
Figure 5: Visualization of attention maps for red-highlighted words: (left) when the model generates the caption in a single pass, and (right) when it performs one more glance with the initial caption.
@@
Figure 6: Reconstructed images from CLIP visual encoder embeddings. We compare results using last-layer features (LF) and multi-layer features (MF). For MF, we extract representations from layers 13, 18, and 23 of the 24-layer ViT-L encoder.
@@
Despite its impressive performance, our specialist shares similar limitations with other MLLMs [65; 45], occasionally producing incorrect predictions. Prior works have interpreted these errors from two perspectives: (i) the hallucination hypothesis, where LLMs, during the reasoning process, describe non-existent elements in the image [5; 39]; (ii) the visual blindness hypothesis, where the visual encoder generates image embeddings which are unclear or ambiguous [66]. Given that our model primarily consists of parameters dedicated to the vision encoder (e.g., 67% of total parameters reside in this component), we hypothesize that the observed errors are more likely a result of visual blindness rather than hallucination. In the following section, we provide an in-depth analysis of this issue and introduce our proposed solution.

### v2

@@ Abstract
Image captioning is fundamental for applications like video-grounded chatbot systems and navigation robots, yet deploying such models on local devices is challenging due to the high computational demands of multimodal LLMs (MLLMs). To address this, we first build lightweight captioning models using a 125M-parameter language model, 56 times smaller than LLaMA-7B, and evaluate their performance not only on single-sentence but on detailed captioning tasks. We obtain surprising results showing that our model can achieve performance comparable to MLLMs, suggesting its potential to serve as a strong captioning specialist for on-device applications. While promising, our model also exhibits a limitation: like other MLLMs, it suffers from occasional captioning errors. We investigate the underlying causes and observe that the problems stem from ineffective attention mechanisms and limited visual representations. To alleviate them, we develop a novel captioning framework, Sharp-Eyed Refinement, which enhances caption quality by refining coarse descriptions into more precise captions. At its core, DeepLens improves visual grounding by re-examining the informative regions identified in the initial glance. Experimental results demonstrate the superiority of our model over both recent lightweight captioning methods and MLLMs in detailed captioning and even in long-range video QA tasks.
Figure 1: (a) Image captioning serves as the cornerstone of diverse applications, such as assistive navigation systems. However, existing captioning models are often too computationally heavy for deployment or suffer from limited performance. (b) The natural process of image description is illustrated: humans first take in the overall scene, then glance at specific regions to notice finer details. (c) Our Sharp-Eyed Refinement framework mimics the human tendency to take a second, more focused glance, allowing the captioning specialist to revise and improve its captioning outputs.
@@ 1 Introduction
Recent progress in image captioning has been driven by the emergence of MLLMs llava ; li2023blip2 , which exploit the advanced language processing capabilities of LLMs zhao2023survey . Building on these advances, image captioning has become a crucial component in various applications, as shown in Figure 1 (a). For example, video-based chatbot systems utilize frame-wise caption generation for temporal understanding wang2023internvid ; moment-gpt ; videotree , while navigation robots construct graph-structured scene descriptions to operate in complex environments xie2024embodied ; jiang2024roboexp .
@@
Despite recent progress, deploying such technologies on local devices remains a significant challenge due to the large computational demands of MLLMs jin2024efficient . Motivated by this, we explore lightweight captioning by implementing a captioning model using OPT-125M zhang2022opt , a model 56 times smaller than LLaMA-7B touvron2023llama in the LLaVA framework llava . To assess its effectiveness, we evaluate it on both single-sentence and detailed captioning tasks, and find that our lightweight model achieves impressive performance. In fact, it not only surpasses previous lightweight captioning models ramos2023smallcap ; huang2023tag2text ; vipcap but also achieves performance comparable to MLLMs, such as LLaVA-7B llava and Instruct-BLIP dai2023instructblip . Taken together, these findings reveal that the complex reasoning capabilities of LLMs are less critical for image captioning tasks that primarily involve enumerating factual details.
@@
Despite these remarkable results, our lightweight model is not without limitations. It occasionally generates inaccurate captions, an issue widely studied in MLLMs, such as LLM’s hallucination bai2024hallucination ; an2024agla and the blindness of vision encoders cambrian ; eyeswideshut . Since the majority of our model’s parameters are dedicated to the visual encoder, we focus on visual blindness as a potential root cause. This issue means that since the vision encoder manipulates visual representation while perceiving the image with its eyes wide shut eyeswideshut , the LLM often struggles to generate accurate outputs.
@@
We take a step towards analyzing the limitations of current captioning models. First, we inspect the model’s attention during a single pass captioning and find that the model struggles to focus on critical regions of the image, instead spreading its attention across the entire image. Moreover, visual representations from a vision encoder like CLIP CLIP lack sufficient detail for accurate caption generation. Specifically, when reconstructing images using these features, significant deviations from the original images are observed. These results highlight the need for more focused and detailed visual representations to provide the LM with clear detection information.
@@
To address these limitations, we propose a novel captioning framework, Sharp-Eyed Refinement, as illustrated in Figure 1. Our method mirrors the human process of image description, where an initial broad understanding of the image is followed by a more focused and detailed analysis. In our framework, our model generates a rough caption, then refines its description. At the heart of our framework is our multimodal connector, DeepLens, which facilitates efficient re-examination of both the image and the initial caption during refinement. The effectiveness of our framework is demonstrated through experiments on single-sentence, detailed captioning, and long-range video QA; representative results are summarized in Figure 2.
@@
Our contributions are summarized as follows: (i) We investigate the practically important yet underexplored topic of lightweight captioning (Section 2). We study lightweight models and demonstrate their potential as strong captioning baselines (Section 3). (ii) Inspired by the human mechanism of ‘one more glance with sharp eyes,’ we introduce a novel refinement framework (Section 4, 5). (iii) We conduct extensive experiments to validate the effectiveness and efficiency of our proposed framework (Section 6). We hope these efforts will inspire further development of efficient captioning models.
@@ 2 Motivation & Scope
Image Captioning as Foundational Technology.
Image captioning is the process of transforming visual information from an image into a natural language description mokady2021clipcap ; hu2022scaling . More than just a standalone task, this technology serves as a fundamental component for a wide range of applications, as outlined in Figure 1 (a). To process video-grounded chatbot systems sarto2023positive ; wang2024internvideo2 ; zhao2023learning ; moment-gpt , image captioning is employed to generate descriptions of multiple frames, which are then integrated into LLMs as prompts to facilitate instruction following. Similarly, exploration robots xie2024embodied ; jiang2024roboexp designed for disaster environments, as well as assistive robots for the visually impaired, utilize captioning to encode observed scenes into a structured graph, supporting navigation and user interactions. In this work, we focus on image captioning as a crucial enabler for these applications.
@@
Recent studies aiming to implement such applications commonly utilize open-source MLLMs wang2023internvid or cloud-based APIs (e.g., OpenAI API) xie2024embodied as image captioners. From a practical perspective, however, these approaches present significant challenges, as (1) open-source MLLMs require substantial computational resources that exceed the capabilities of edge devices like smartphones cai2020tinytl ; song2023ecotta , while (2) cloud-based APIs depend on a stable internet connection, which is likely to be unavailable in disaster situations. Table 1 presents the number of parameters used by recent MLLMs and their NVIDIA GPU memory requirements, assuming FP16 precision. Given that these applications require repetitive captioning process across multiple scenes zhang2023simple , the feasibility of current models for real-world deployment remains uncertain. This resource-demand mismatch highlights the need for lightweight captioning models.
@@
Generalist vs. Specialist.
Following previous works ye2024mplug ; mome ; wu2024visionllm , we use the term generalists to denote models trained on diverse data for broad purposes (e.g., current MLLMs llava15 ; llavanext ), and specialists to denote models trained solely on task-specific data (e.g., small captioning models ramos2023smallcap ; vipcap )."
@@ 3 Exploring Lightweight Captioning
To address the challenges outlined above, this section presents our design of a lightweight captioning specialist and provides comprehensive analyses of its effectiveness.
Model construction. To construct a lightweight captioning model, we initially focus on minimizing reliance on LLMs, as they contribute to the majority of computational overhead in MLLMs (e.g., 96% in LLaVA-7B llava is attributed to LLaMA touvron2023llama ). To this end, we build our captioning specialist by replacing the LLaMA-7B used in LLaVA construction with OPT-125M zhang2022opt , a language model 56 times smaller. It is notable that OPT-125M is significantly compact unlike modern small-scale language models exceeding 500M parameters.
Implementation details. We use the publicly available LLaVA-1.5 llava15 code. Other than this modification, all training strategies follow the original code, including batch size, learning rate, and updating both the language model and the connector during fine-tuning while freezing the vision encoder. More details are available in Section I.1 and our source code [^1] https://github.com/junha1125/Lightweight-Captioner . All our experiments are conducted using two NVIDIA A6000 GPUs.
Datasets and Evaluations. The multimodal connector is first pretrained on the Caption Concept-balanced 558K llava15 , after which the model is fine-tuned on task-specific datasets such as COCO mscoco , DCI dci , or ShareGPT4V chen2024sharegpt4v , as introduced later. We consider traditional metrics such as BLEU papineni2002bleu , CIDEr vedantam2015cider , and BERTScore bertscore , as well as MLLM-as-a-Judge mllmjudge ; clair with GPT-4o-mini.
@@ 3.1 Results on Single Sentence Captioning
MS COCO Captions mscoco , the most widely used captioning benchmark, consists of 113k images and 5 captions per image. One caption typically contains a single sentence of about 10 words. The results in Table 2 show that our model outperforms previous small captioning works huang2023tag2text ; wan2024locca that utilize fewer than 1B parameters. Specifically, our model achieves CIDEr scores that are 6.9 higher than SmallCap ramos2023smallcap , which also uses OPT-125M as its LM backbone. Considering that our model does not involve any newly proposed methods, these results show surprisingly strong performance. We further discuss these results in Section E.1. Furthermore, we compare our model against MLLM generalists chen2023pali ; bai2023qwenvl . Despite requiring significantly fewer computational resources, our model achieves comparable performance to these larger counterparts.
MS COCO mscoco
model venue # data # params B@4 papineni2002bleu MET meteor CIDEr vedantam2015cider BERT bertscore CLAIR clair GPT mllmjudge
G InstructBLIP dai2023instructblip NeurIPS23 130M> 8.2B 38.0 29.4 127.8 69.1 - -
Unified-IOXL lu2023unifiedio ICLR23 130M 7.3B 37.0 29.5 123.6 68.2 - -
Shikra chen2023shikra arXiv23 - 7.2B - - 117.5 - - -
@@
ViPCap vipcap AAAI25 4M 225M 37.7 28.6 122.9 - - -
[…]
ShareGPT4V chen2024sharegpt4v & DCI dci
model venue # data # params B@4 papineni2002bleu CIDEr vedantam2015cider BERT bertscore CAPT dong2024benchmarking CLAIR clair GPT mllmjudge
[…]
LLaVA-1.5 llava15 CVPR24 1M 7.3B 10.6 36.1 36.6 48.6 $\pm$ $\pm$
EyesWideShut eyeswideshut CVPR24 1M 7.6B 11.6 36.5 37.2 49.0 - -
[…]
SmallCap* ramos2023smallcap CVPR23 0.5M 450M 14.5 20.1 28.9 23.3 - -
@@
GLaMM rasheed2024glamm
model venue # data # params B@4 papineni2002bleu CIDEr vedantam2015cider BERT bertscore CAPT dong2024benchmarking CLAIR clair GPT mllmjudge
@@
Table 2: Comparison of captioning performance. We evaluate our specialist and existing captioning models. Despite not introducing any newly proposed methods, our model achieves strong performance. Here, ‘G’ represents a MLLM generalist, while ‘S’ denotes the small captioning model. ‘*’ indicates models that are fine-tuned in this study, as they were not trained for detailed captioning tasks.
Figure 3: Qualitative results on detailed captioning. Despite its small size, our specialist generates effective image captions, with additional examples shown in Figure 15. For comparison, results from generalist models are provided in Figure 12 and Figure 13.
@@ 3.2 Results on Detailed Captioning
For this experiment, we fine-tune our model on a combination of the DCI dci and ShareGPT4V chen2024sharegpt4v datasets, and additionally utilize the GLaMM rasheed2024glamm dataset. We then evaluate the model on the test splits of these datasets. Further details are provided in Section G.2.
@@
We initially anticipated that our specialist would struggle with this challenging task, given that OPT-125M possesses limited language processing capabilities, presumably due to its small parameter count, as shown in Table 4. However, the results in Table 2 and Figure 3 reveal that our model performs surprisingly well on detailed captioning tasks, contrary to our expectations. It should be noted that the vision encoder remains frozen while only OPT-125M is updated, meaning that detailed caption generation is largely handled by OPT-125M, with the vision encoder offering visual cues.
Figure 4: Generation examples from OPT-125M. Surprisingly, despite its limited capacity, our lightweight specialist performs image captioning effectively.
Results Summary.
@@
Although MLLMs rely on the complex reasoning abilities of LLMs for tasks such as VQA and instruction following, our experiments indicate that captioning—primarily involving the enumeration of factual objects—does not significantly require such capabilities. This observation suggests that a lightweight captioning model can serve as an efficient and practical alternative for captioning-based applications.
@@
Note 1: Factual image captioning does not significantly require the complex reasoning abilities of LLMs and can be handled by a lightweight language model.
@@
Figure 5: Visualization of attention maps for highlighted words. The left figure shows attention during single-pass caption generation, while the right figure illustrates the refinement process with one more glance guided by the initial caption.
@@
Figure 6: Image reconstruction using visual representations from CLIP CLIP . Results are shown for last-layer features (i.e., LF) and multi-layer features (i.e., MF), where MF corresponds to representations extracted from layers 13, 18, and 23 of the 24-layer OpenAI CLIP ViT-L/14@336.
@@
Despite its impressive performance, our model shares similar limitations with other MLLMs cambrian ; lu2024deepseekvl , occasionally producing incorrect predictions. Prior works have interpreted these errors from two perspectives: (i) the hallucination hypothesis, where LLMs, during response generation, describe non-existent elements in the image bai2024hallucination ; li2023evaluatingobjecthallucinationlarge ; (ii) the visual blindness hypothesis, where the visual encoder generates visual representations which are unclear or ambiguous eyeswideshut . Given that our model primarily consists of parameters dedicated to the vision encoder (e.g., 67% of total parameters), we hypothesize that the observed errors are more likely a result of visual blindness. In the following section, we provide an in-depth analysis of this issue and introduce our proposed solution.